OpenStudy (jazzyluvsu123):

A copy shop leases a copy machine for $256.25 per month plus $0.008per copy made on the machine. If the copy shop makes 12,480 copies in a month, what will be the lease payment? Help me work this out FOR A FREE MEDAL @agent0smith

OpenStudy (anonymous):

okay so we are going to declare a variable. X= the # of copies then we will set up an equation y=.008x+256.25. So if you are making 12,480 copies then you would substitute in parentheses that number for x and solve. :)

OpenStudy (jazzyluvsu123):

so .008 + 256.25 = my answer right ?

OpenStudy (anonymous):

no it would be (.008)(12480) + 256.25 because if x = the number of copies you would put the number of copies given to you in the question in for x.

OpenStudy (jazzyluvsu123):

im in 7th grade and I don't under stand if I am dividing, *, +,Or -

OpenStudy (jazzyluvsu123):

@jdifatta

OpenStudy (anonymous):

No problem, but do you understand what you need to do.

OpenStudy (jazzyluvsu123):

no sorry not really can you like help me through the steps imnot really good at math

OpenStudy (jazzyluvsu123):

i think i got it now is it 12,480 x .008 = something + 256.25 = my answer or am i wrong

OpenStudy (jazzyluvsu123):

@jdifatta

OpenStudy (agent0smith):

$0.008*12,480 + $256.25 = ?

OpenStudy (jazzyluvsu123):

is that how i do it ?

OpenStudy (anonymous):

Well, if they make 12480 copies.. and they have to pay 0.008 per copy... 12480 x 0.008 is how much they have to pay for all the copies And then add the base monthly lease payment of 256.25 Does that make sense? :)

OpenStudy (anonymous):

yes let me draw it up for you

OpenStudy (jazzyluvsu123):

yes let work it out and I will post the answer and tell me if i am right ?! ok

OpenStudy (anonymous):